Density of 2.05M solution of acetic acid in water is 1.02g/ml. find molality of same solution.

Grade:12th pass

2 Answers

Anurag Sen
6 years ago
Weight of acetic acid =2.05×60=1232.05×60=123Weight of the solution =1000×1.02=10201000×1.02=1020∴∴ Weight of water=(1020−123)=897g(1020−123)=897g∴∴ Molality =2.05×1000897=2.285
Anurag Sen
6 years ago
Weight of acetic acid =2.05×60=1232.05×60=123Weight of the solution =1000×1.02=10201000×1.02=1020∴∴ Weight of water=(1020−123)=897g(1020−123)=897g∴∴ Molality =2.05×1000897=2.285 that`s it